What is the present value of the following cash flow stream? 

Year 01 to year 10: $300 

Year 11 to year 20: $200 

From year 21 and onwards: S100 


Assume a 10% discount rate.
Answer #1

There are two parts of this question:

  1. Find the PV of the CFs for years 1-20
  2. Find the PV value of the perpetuity for years 21 onward

Then we sum up the two PVs

100 Perpetuity = = 1000

Now we find the PV of the perpetuity and the CFs from years 1-20

We should know that the perpetuity is at time 20 and not time 21

So the PV is :

Year CF Discount Factor Discounted CF
0 $              -   1/(1+0.1)^0= 1 1*0=                -  
1 $     300.00 1/(1+0.1)^1= 0.909090909 0.909090909090909*300=       272.73
2 $     300.00 1/(1+0.1)^2= 0.826446281 0.826446280991735*300=       247.93
3 $     300.00 1/(1+0.1)^3= 0.751314801 0.751314800901578*300=       225.39
4 $     300.00 1/(1+0.1)^4= 0.683013455 0.683013455365071*300=       204.90
5 $     300.00 1/(1+0.1)^5= 0.620921323 0.620921323059155*300=       186.28
6 $     300.00 1/(1+0.1)^6= 0.56447393 0.564473930053777*300=       169.34
7 $     300.00 1/(1+0.1)^7= 0.513158118 0.513158118230706*300=       153.95
8 $     300.00 1/(1+0.1)^8= 0.46650738 0.466507380209733*300=       139.95
9 $     300.00 1/(1+0.1)^9= 0.424097618 0.424097618372485*300=       127.23
10 $     300.00 1/(1+0.1)^10= 0.385543289 0.385543289429531*300=       115.66
11 $     200.00 1/(1+0.1)^11= 0.350493899 0.350493899481392*200=         70.10
12 $     200.00 1/(1+0.1)^12= 0.318630818 0.318630817710357*200=         63.73
13 $     200.00 1/(1+0.1)^13= 0.28966438 0.289664379736688*200=         57.93
14 $     200.00 1/(1+0.1)^14= 0.263331254 0.26333125430608*200=         52.67
15 $     200.00 1/(1+0.1)^15= 0.239392049 0.239392049369163*200=         47.88
16 $     200.00 1/(1+0.1)^16= 0.217629136 0.217629135790149*200=         43.53
17 $     200.00 1/(1+0.1)^17= 0.197844669 0.197844668900135*200=         39.57
18 $     200.00 1/(1+0.1)^18= 0.17985879 0.179858789909214*200=         35.97
19 $     200.00 1/(1+0.1)^19= 0.163507991 0.163507990826558*200=         32.70
20 $     200.00 1/(1+0.1)^20= 0.148643628 0.148643628024143*200=         29.73
20 $ 1,000.00 1/(1+0.1)^20= 0.148643628 0.148643628024143*1000=       148.64
NPV = Sum of all Discounted CF 2,465.81